# DocETL: Agentic Query Rewriting and Evaluation for Complex Document Processing

### v2

@@ Abstract.
We present DocETL, a system that optimizes complex document processing pipelines, while accounting for LLM shortcomings. DocETL offers a declarative interface for users to define such pipelines and uses an agent-based approach to automatically optimize them, leveraging novel agent-based rewrites (that we call rewrite directives), as well as an optimization and evaluation framework. We introduce (i) logical rewriting of pipelines, tailored for LLM-based tasks, (ii) an agent-guided plan evaluation mechanism that synthesizes and orchestrates task-specific validation prompts, and (iii) an optimization algorithm that efficiently finds promising plans, considering the latencies of agent-based plan generation and evaluation. Our evaluation on four different unstructured document analysis tasks demonstrates that DocETL finds plans with outputs that are $25$ to $80\%$ more accurate than well-engineered baselines, addressing a critical gap in unstructured data analysis. DocETL is open-source at docetl.org, and as of November 2024, has amassed over 1.3k GitHub Stars, with users spanning a variety of domains.
@@ 1. Introduction
DocETL is open-source and available on GitHub [^2] https://github.com/ucbepic/docetl. As of November 2024, it has already amassed 1.3k+ GitHub stars, and has been used for pipelines ranging from domain-specific analysis (e.g., legal, climate science) to enterprise and personal productivity (e.g., analyzing customer support tickets, emails); over 300 users have joined the corresponding Discord server.
@@
We describe DocETL’s programming model and operators in Section 2; our new LLM-centric rewrite directives in Section 3, the agentic optimizer that applies them, and evaluates the resulting plans, as well as the overall framework for optimization in Section 4. We present our initial evaluation in Section 5, where we demonstrate that across four unstructured document analysis tasks, DocETL finds plans that are 25 to 80% more accurate than baselines. We then reflect on next steps in Section 6, and discuss related work in Section 7.
@@ 2.1. Programming Model
DocETL processes collections of documents. A document comprises a set (or dictionary) of key (or equivalently, attribute)-value pairs, represented as a JSON object. For example, a police record could be a set of key-value pairs, where one key corresponds to the OCR output of the PDF, while other keys could capture metadata such as responding agency, file name, or creation date. A collection of documents or dataset, is a JSON array. This data representation lets us flexibly handle various data types and degrees of structure and easily reference data within operation prompts. Documents can be nested, e.g., a police record may contain an array of related_documents that each contain witness statements or evidence logs that are further nested.
@@ 2.2.1. Map
Extract any instances of officer misconduct or procedural violations. For each instance, provide:
1. The name of the officer involved
2. A brief description of the misconduct or violation
@@ 2.2.3. Resolve
Compare the following two officers from police records:
Officer 1: {{ input1.officer_name }} mentioned in: {{ input1.record_txt }} and Officer 2: {{ input2.officer_name }} mentioned in: {{ input2.record_txt }}
Are these names likely referring to the same officer?
@@ 2.3.3. Gather
This particular configuration includes the full content of the document’s first chunk, summaries of intermediate chunks, and the current chunk itself. Figure 3 demonstrates different ways to render chunks. The gather operation is highly flexible in rendering contextual information, allowing for the inclusion of full chunks (as in (ii)), portions of chunks (as in (i)), or transformations (e.g., summaries) of chunks (as in (iii)). Importantly, there may be map operations between the split and gather steps—allowing for the generation of additional context (such as summaries) that can be used to augment each chunk, before downstream processing.
The output adds a new attribute to each input document, containing the rendered chunk with its peripheral context, with with special tags that demarcate what is the chunk and what is peripheral context. This approach ensures that each chunk is processed with necessary context, maintaining the coherence and structure of the original document even when split across multiple chunks. For additional details, see Appendix A.
@@ 3. Rewrite Directives
Throughout this section, we adopt the following notation: given operators $A$ and $B$, we denote their composition as $A\to B$, where $(A\to B)(D)=B(A(D))$. For independent execution of operators, we use $A\parallel B$ to indicate that $A$ and $B$ are executed on the same input, independent of each other. For readability, we may drop arguments—e.g., $\text{Map}_{x}(D)$ becomes $\text{Map}_{x}$. Similarly, we omit subscripts except when the same operator appears in multiple places. We further refer to the text content of the document, usually stored as one of the attributes, interchangeably with the document itself, for simplicity.
@@ 3.1.1. Document Chunking (Map)
Ignoring the blue annotations, this directive rewrites map to: split the document into multiple chunks, gather peripheral context for each chunk, apply a modified map operation per chunk, and reduce the results. The prompt for $\text{Map}_{y}$ may explicitly state that only a portion of the original document is being processed.
To provide more flexibility and optimization opportunities, we introduce smaller decomposition directives, for steps (2)–(5) above:
@@
  • Header Lineage Context and Summarization (3): This directive introduces two parallel map operations: $\text{Map}_{h}$ for extracting hierarchical information (e.g., headers), and $\text{Map}_{s}$ for generating summaries of chunks. This allows us to provide each chunk with its relevant hierarchical context (e.g., parent headers for headers in a chunk) and/or a summary of preceding content.
@@ 3.3. Projection Synthesis
  • Isolating (11): For map operations with independent subtasks, this directive splits them into separate projections to run in parallel, followed by a reduce step. For instance, customer feedback analysis could involve parallel projections to classify sentiment, identify features, and flag urgent issues.
@@
Overall, our rewrite directives reflect our key insight: in complex document processing tasks, it is impossible to determine an optimal pipeline given the infinite search space, difficulty, and ambiguity. Rewrite directives provide a scaffold for systematically exploring this space, especially when coupled with opportunistic decomposition into problematic operations (as will be described in subsequent sections). The effectiveness of specific directives varies by context and is hard to predict—we plan to study their relative utility in future work. Finally, as a byproduct of this search process to find sufficiently accurate pipelines, we obtain interpretable pipelines, since the operators use natural language prompts.
@@ 4.1. Optimization Approach
DocETL employs a top-down optimization approach that considers both individual operations and sub-pipelines, as outlined in Algorithm 1 and visualized in Figure 1. Intuitively, we move from left to right, opting (recursively) to decompose any operations for which the accuracy is inadequate (as determined by the LLM validators). The process can be summarized as follows:
@@
    Rewrite Directive Application and Recursive Optimization: When optimization is needed, we apply matching rewrite directives to the sub-pipeline or individual operation. As illustrated in Figure 1, we explore various rewrite directives from Section 3. For each applicable directive, an LLM agent synthesizes new operations and configurations (e.g., prompts, output schemas) to match the directive. When new operations are created, we immediately optimize them, recursively, before continuing with the current optimization, as shown by the nested “Apply Rewrites” rectangles in the figure. This opportunistic approach allows us to explore more refined plans efficiently (Algorithm 2, lines 10-11).
@@
    Pipeline Update: We integrate the selected optimized plan into the overall pipeline, replacing the original operation or sub-pipeline (Algorithm 1, lines 9-12).
@@ 4.2.1. Generation Agents
Chunk Sizes. Our chunking approach explores sizes ranging from 20% to 80% of the LLM’s context limit. For each chunk size, we generate a set of gather configurations to retain relevant context from surrounding chunks. The creation of these gather configurations is based on the ratio of chunk size to document size.
@@
Batch Sizes. For reduce operations, optimal batch sizes (i.e., the number of documents aggregated at once, in a single prompt) are not obvious and require experimentation. Our agent tests sizes from 20% to 100% of the maximum input fitting the LLM’s context window, generating and evaluating multiple fold prompts for each. Our evaluations reveal task-dependent optimal batch sizes, highlighting the need for further research in this area—some tasks perform best with the smallest batch size (e.g., extracting distinct names), while others peak at a middle batch size, as shown in Section 5.
@@ 5. Evaluation
Overall, we find that DocETL’s plans yield 25% to 55% improvements in task-specific accuracy metrics such as precision, recall, and F1 score. We first consider three complex document processing tasks: legal contract analysis, declassified article analysis, and video game review analysis. These tasks represent different challenges: extracting structured information embedded within the semantic content of unstructed data, resolving entities and summarizing their information across documents, and reasoning about temporal consistency across long documents. For the first task, we compare against LOTUS (Patel et al., 2024) and Palimpzest (Liu et al., 2024b), both recent preprints. We exclude these systems from the second and third tasks as they lack support for entity resolution and long documents (i.e., resolve, split, gather). We additionally evaluate DocETL on the challenging Biodex text classification (D’Oosterlinck et al., 2023) from the LOTUS paper, their only task with sub-70% accuracy, thereby benefiting from accuracy optimization. On this task, DocETL’s optimized pipeline achieves 33 to 80% improvements in rank precision compared to LOTUS. Finally, we report on a case study of applying DocETL to police misconduct, as well as anecdotes from open-source users.
For all pipelines, we use the gpt-4o-mini model from OpenAI, and we run the experiments on a 2021 Macbook Pro with an M1 chip. The DocETL optimizer uses gpt-4o-mini, except in the police misconduct task in Section 5.5, where we use gpt-4o.
@@ 5.1. Legal Contract Analysis
We evaluate on a subset, the first 50 contracts, comparing extracted clauses against ground truth. An extraction is considered correct if (i) the clause type matches the ground truth, and (ii) the extracted text span has a Jaccard similarity greater than a specific threshold (set as $>0.15$) with the ground truth span. This similarity threshold accommodates variation in LLM outputs while ensuring the model has correctly identified the clause’s location; it is set fairly low because we provide no training examples, so the LLM does not know how little or how much to extract—but large enough to ensure some match. We set other values for this and found the comparisons to be similar. We measure precision, recall, and F1.
@@ 5.1.1. Implementations
We have three baselines:
@@ 4. Effective Date: The date when the contract is effective
    DocETL’s Optimized Plan: DocETL’s optimizer transforms the single map operation into a parallel decomposition with 21 map operations, each extracting 1-3 semantically related spans (e.g., grouping agreement and effective date extractions), followed by a reduce to combine all extracted clauses. Notably, the optimizer chose parallel decomposition (directive 11) over document chunking, suggesting that LLMs excel at focused extraction of small amounts of information even from lengthy documents.
@@ 5.1.2. Results
The results are shown in Table 2. DocETL’s optimized plan performs significantly better than all baselines, achieving a 25.1% improvement in F1 and 54.5% improvement in recall over LOTUS, the next best plan. LOTUS and the unoptimized DocETL pipeline achieve similar scores, as neither method employs any accuracy optimization. Interestingly, Palimpzest’s optimizer selected a code-based plan rather than an LLM-based one for this task, despite selecting the maxQuality plan. Palimpzest’s lower performance on this specific task may be due to difficulties in configuring its schema-only approach.
While the optimized pipeline’s cost and runtime are higher (Table 3), we prioritize accuracy, which often requires increased computational costs. The higher runtime and cost stems from the increased number of LLM calls in the parallel map operations, plus an additional reduce operation to combine their results. Further parallelism could help reduce the runtimes further, but this is not our focus. Costs will decrease as LLM pricing continues to fall—they have fallen by 1000$\times$ in 3 years, with a predicted drop of 10$\times$ per year (Appenzeller, 2024)—and they become negligible when using open-source models. The optimization cost is only \$1.58 (using gpt-4o-mini for the optimizer’s agents) and does not increase with dataset size, as it is done on a sample.
Table 2. Legal Contract Analysis Results
System Avg Precision Avg Recall Avg F1
DocETL (Unopt.) 0.305 0.451 0.364
DocETL (Opt.) 0.394 0.731 0.474
LOTUS 0.350 0.473 0.379
Palimpzest 0.059 0.013 0.022
@@
DocETL (Unopt.) 23.43 0.10 N/A
DocETL (Opt.) 180.30 1.70 1.58
@@
Palimpzest uses DSPy (Khattab et al., 2024) which makes LLM calls opaque, preventing cost tracking.

### v3

@@ Abstract.
We present DocETL, a system that optimizes complex document processing pipelines, while accounting for LLM shortcomings. DocETL offers a declarative interface for users to define such pipelines and uses an agent-based approach to automatically optimize them, leveraging novel agent-based rewrites (that we call rewrite directives), as well as an optimization and evaluation framework. We introduce (i) logical rewriting of pipelines, tailored for LLM-based tasks, (ii) an agent-guided plan evaluation mechanism that synthesizes and orchestrates task-specific validation prompts, and (iii) an optimization algorithm that efficiently finds promising plans, considering the latencies of agent-based plan generation and evaluation. Our evaluation on four different unstructured document analysis tasks demonstrates that DocETL finds plans with outputs that are $21$ to $80\%$ more accurate than well-engineered baselines. DocETL is open-source at docetl.org, and as of March 2025, has amassed over 1.7k GitHub Stars, with users spanning a variety of domains.
@@ 1. Introduction
DocETL is open-source and available on GitHub [^2] https://github.com/ucbepic/docetl. As of March 2025, it has already amassed 1.7k+ GitHub stars, and has been used for pipelines ranging from domain-specific analysis (e.g., legal, climate science) to enterprise and personal productivity (e.g., analyzing customer support tickets, emails); over 400 users have joined the corresponding Discord server.
@@
We describe DocETL’s programming model and operators in Section 2; our new LLM-centric rewrite directives in Section 3, the agentic optimizer that applies them, and evaluates the resulting plans, as well as the overall framework for optimization in Section 4. We present our initial evaluation in Section 5, where we demonstrate that across four unstructured document analysis tasks, DocETL finds plans that are 21 to 80% more accurate than baselines. We then reflect on next steps in Section 6, and discuss related work in Section 7.
@@ 2.1. Programming Model
DocETL processes collections of documents. A document comprises a set (or dictionary) of key (or equivalently, attribute)-value pairs, represented as a JSON object. For example, a police record could be a set of key-value pairs, where one key corresponds to the OCR output of the PDF, while other keys could capture metadata such as agency, file name, or creation date. A collection of documents or dataset, is a JSON array. This data representation lets us handle various data types and degrees of structure and easily reference data within operation prompts. Documents can be nested, e.g., a police record may contain an array of related_documents that each contain witness statements or evidence logs that are further nested.
@@ 2.2.1. Map
Extract any instances of officer misconduct or procedural violations. For each instance, provide the name of the officer involved and a brief description of the misconduct or violation.
@@ 2.2.3. Resolve
Compare the following two officers from police records. Officer 1: {{ input1.officer_name }} mentioned in: {{ input1.record_txt }} and Officer 2: {{ input2.officer_name }} mentioned in: {{ input2.record_txt }} Are these names likely referring to the same officer?
@@ 2.3.3. Gather
This particular configuration includes the full content of the document’s first chunk, summaries of intermediate chunks, and the current chunk itself. Figure 3 demonstrates different ways to render chunks. The gather operation is highly flexible in rendering contextual information, allowing for the inclusion of full chunks (as in (ii)), portions of chunks (as in (i)), or transformations (e.g., summaries) of chunks (as in (iii)). Importantly, there may be map operations between the split and gather steps—allowing for the generation of additional context (such as summaries) that can be used to augment each chunk, before downstream processing. The output adds a new attribute to each input document, containing the rendered chunk with its peripheral context, with with special tags that demarcate what is the chunk and what is peripheral context. For additional details, see Appendix A.
Overall, in designing the DocETL DSL, we unified various single-document transformations (e.g., extraction, summarization) under map and filter operators, letting users express intent through prompts rather than learning multiple specialized operators. But for cross-document operations, we created distinct operators that capture specific processing patterns. For example, while resolve could theoretically be implemented using equijoin, reduce, and another equijoin, having a dedicated operator allows us to know that the user’s intent is actually entity resolution, so we can better optimize the pipeline. Additionally, we distinguish gather from reduce because they serve different purposes: reduce performs many-to-one aggregation, whereas gather preserves cardinality while enriching documents with context—similar to SQL windowing functions.
@@ 3. Rewrite Directives
Throughout this section, we adopt the following notation: given operators $A$ and $B$, we denote their composition as $A\to B$, where $(A\to B)(D)=B(A(D))$. For independent execution of operators, we use $A\parallel B$ to indicate that $A$ and $B$ are executed on the same input, independently. For readability, we may drop arguments—e.g., $\text{Map}_{x}(D)$ becomes $\text{Map}_{x}$. Similarly, we omit subscripts except when the same operator appears in multiple places. We further refer to the text content of the document, usually stored as one of the attributes, interchangeably with the document itself, for simplicity. The arrow $\Rightarrow$ denotes a (semantic) rewrite of the operator (or operator sequence) on the left into the form on the right.
@@ 3.1.1. Document Chunking (Map)
Ignoring the purple annotations, this directive rewrites map to: split the document into multiple chunks, gather peripheral context for each chunk, apply a modified map operation per chunk, and reduce the results. The prompt for $\text{Map}_{y}$ may explicitly state that only a portion of the original document is being processed. To provide more flexibility and optimization opportunities, we introduce smaller decomposition directives, for steps (2)–(5) above:
@@
  • Header Lineage Context and Summarization (3): This directive introduces two independent map operations: $\text{Map}_{h}$ for extracting hierarchical information (e.g., headers), and $\text{Map}_{s}$ for generating summaries of chunks. This allows us to provide each chunk with its relevant hierarchical context (e.g., parent headers for headers in a chunk) and/or a summary of preceding content.
@@ 3.3. Projection Synthesis
  • Isolating (11): For map operations with independent subtasks, this directive splits them into separate projections to run in parallel, followed by a reduce step. For instance, customer feedback analysis could involve isolated projections to classify sentiment, identify features, and flag urgent issues.
@@
Overall, our rewrite directives reflect our key insight: in complex document processing tasks, it is impossible to determine an optimal pipeline given the infinite search space, difficulty, and ambiguity. Rewrite directives provide a scaffold for systematically exploring this space, especially when coupled with opportunistic decomposition into problematic operations (as will be described in subsequent sections). The effectiveness of specific directives varies by context and is hard to predict. Finally, as a byproduct of this search process to find sufficiently accurate pipelines, we obtain interpretable pipelines, since the operators use natural language prompts.
@@ 4.1. Optimization Approach
DocETL employs a top-down optimization approach that considers both individual operations and sub-pipelines, as outlined in Algorithm 1 and visualized in Figure 1. We move from left to right, opting (recursively) to decompose any operations for which the accuracy is inadequate (as determined by the LLM validators). We summarize the process:
@@
    Rewrite Directive Application and Recursive Optimization: When optimization is needed, we apply matching rewrite directives to the sub-pipeline or individual operation. As illustrated in Figure 1, we explore rewrite directives from Section 3. For each applicable directive, an LLM agent synthesizes new operations and configurations (e.g., prompts, output schemas) to match the directive. On the creation of a new operation, we immediately optimize it, recursively, before continuing with the current optimization, as shown by the nested “Apply Rewrites” rectangles in the figure. This opportunistic approach allows us to explore more refined plans efficiently (Algorithm 2, lines 10-11).
@@
    Pipeline Update: We integrate the selected optimized plan into the pipeline, replacing the original operation or sub-pipeline (Algorithm 1, lines 9-12).
@@ 4.2.1. Generation Agents
Chunk Sizes. Our chunking approach explores five sizes ranging from 15% to 75% of the LLM’s context limit, uniformly sampled. We also explore chunk sizes based on percentages of the average document length; similarly six sizes ranging from 15% to 100%, uniformly sampled. also For each chunk size, we generate a set of gather configurations to retain relevant context from surrounding chunks. The creation of these gather configurations is based on the ratio of chunk size to document size.
@@
Batch Sizes. For reduce operations, optimal batch sizes (i.e., the number of documents aggregated at once, in a single prompt) are not obvious and require experimentation. Our agent tests sizes from 20%, 40%, 60%, 75%, to 90% of the maximum input fitting the LLM’s context window, generating and evaluating multiple fold prompts for each. Our evaluations reveal task-dependent optimal batch sizes, highlighting the need for further research in this area—some tasks perform best with the smallest batch size (e.g., extracting distinct names), while others peak at a middle batch size, as shown in Section 5.
@@ 5. Evaluation
Overall, we find that DocETL’s plans yield 21% to 80% improvements in task-specific accuracy metrics such as precision, recall, and F1 score. We first consider three complex document processing tasks: legal contract analysis, declassified article analysis, and video game review analysis (Sections 5.1, 5.3 and 5.2). These tasks represent different challenges: extracting structured information embedded within the semantic content of unstructured data, resolving entities and summarizing their information across documents, and reasoning about temporal consistency across long documents. For the legal contract analysis (Section 5.1), we compare against both recent LLM-powered systems (LOTUS (Patel et al., 2024), Palimpzest (Liu et al., 2024b), and Aryn (Anderson et al., 2024)) and traditional NLP baselines using spaCy (Honnibal et al., 2020) or NLTK (Bird et al., 2009). For the video game review (Section 5.2) and declassified article (Section 5.3) tasks, we compare only against non-LLM baselines, as LOTUS, Palimpzest, and Aryn lack support for entity resolution and documents exceeding LLM context windows. For each task, our evaluation includes both task-specific metrics (customized variations of precision and recall) as well as a hallucination rate to measure factual consistency. Then, we evaluate DocETL on the challenging Biodex text classification task (LOTUS’s only task with sub-70% accuracy), where our optimized pipeline achieves 33 to 80% improvements in rank precision (Section 5.4). We conclude with case studies examining DocETL’s application in real-world police misconduct identification, the effectiveness of LLM agent rewrites, and insights from user adoption (Sections 5.5, 5.6 and 5.7).
For all pipelines, we use the gpt-4o-mini model from OpenAI, and we run the experiments on a 2021 Macbook Pro with an M1 chip. The DocETL optimizer uses gpt-4o-mini, except in the Biodex task in Section 5.4, where we use gpt-4o.
@@ 5.1. Legal Contract Analysis
We evaluate on the first 50 contracts, comparing extractions against ground truth. An extraction is considered correct if (i) the clause type matches, and (ii) the extracted text span’s Jaccard similarity with the ground truth span $>0.15$. This threshold accommodates variation in LLM outputs while ensuring the model has correctly identified the clause’s location; it is set fairly low because we provide no training examples, so the LLM does not know how much to extract—but large enough to ensure some match. We set other values for this and found the comparisons to be similar. We measure precision, recall, F1, and hallucination rate (proportion of extracted clauses not matching our 41 predefined categories).
@@ 5.1.1. Implementations
We have five baselines:
@@ 4. Effective Date: The date when the contract is effective
    Non-LLM Baseline: We write a program, using the spaCy library (Honnibal et al., 2020), to loop over all clause types and extract the most semantically similar sentence (above a threshold of 0.9). We use spaCy’s sentence splitter and embedding model, tok2vec.
  5. (5)
    Aryn Baseline: We implement extraction using Aryn’s llm_query operation with the same prompt as our LOTUS baseline, and the same output normalization procedure used with LOTUS to handle parsing errors and format inconsistencies.
  6. (6)
    DocETL’s Optimized Plan: DocETL’s optimizer transforms the single map operation into an isolated projection decomposition with 21 independent map operations, each extracting 1-3 semantically related spans (e.g., grouping agreement and effective date extractions), followed by a reduce to combine all extracted clauses. Notably, the optimizer chose isolated projection (directive 11) over document chunking, suggesting that LLMs excel at focused extraction of small amounts of information even from lengthy documents.
@@ 5.1.2. Results
The results are shown in Table 2. DocETL’s optimized plan performs significantly better than all baselines, achieving a 21.4% improvement in F1 over LOTUS, the next best LLM-based plan, and a 67% improvement in recall over the unoptimized DocETL pipeline—with no hallucinations. LOTUS, Aryn, and the unoptimized DocETL pipelines achieve similar scores and hallucination rates (6.9-7.3%). The non-LLM baseline achives much lower scores than the LLM-based methods, as well as longer text spans—because spans are forced to be at sentence-level granularity, which could be longer than necessary for short clauses like “document name” or “agreement date.” Interestingly, Palimpzest’s optimizer selected a code-based plan rather than an LLM-based one for this task—perhaps explaining its lower score. Palimpzest’s lower performance on this specific task may be due to difficulties in configuring its schema-only approach.
While the optimized pipeline’s cost and runtime are higher (Table 3), we prioritize accuracy, which often requires increased computational costs. The higher runtime and cost stems from the increased number of LLM calls in the new map operations, plus an additional reduce operation to combine their results. Further parallelism could help reduce the runtimes further, but this is not our focus. Costs will decrease as LLM pricing continues to fall—they have fallen by 1000$\times$ in 3 years, with a predicted drop of 10$\times$ per year (Appenzeller, 2024)—and they become negligible when using open-source models. The optimization cost is only \$1.58 (using gpt-4o-mini for the optimizer’s agents) and does not increase with dataset size, as it is done on a sample.
Table 2. Legal Contract Analysis Results.
System Avg Precision Avg Recall Avg F1 Avg # Chars Avg Hallucination Rate
DocETL (Opt.) 0.401 0.719 0.477 162.60 0.000
DocETL (Unopt.) 0.341 0.430 0.379 49.35 0.072
LOTUS 0.402 0.471 0.393 46.301 0.073
Palimpzest 0.059 0.013 0.022 35.10 0.000
Aryn 0.450 0.370 0.352 49.56 0.069
Non-LLM 0.224 0.219 0.190 212.73 0.000
@@
DocETL (Opt.) 180.30 1.46 1.58
DocETL (Unopt.) 23.43 0.08 N/A
@@
Aryn 52.53 Unknown N/A
Non-LLM 217.99 0.00 N/A
Costs are not reported by the system.